A triangular garden has an area of 90 sq feet. It's height is 8 feet more than its base. Find the measure of the base.
Answer by nerdybill(7384)

A triangular garden has an area of 90 sq feet. It's height is 8 feet more than its base. Find the measure of the base.
.
Since:
A = (1/2)bh
where
A is area
b is base
h is height
.
90 = (1/2)b(b+8)
180 = b(b+8)
180 = b^2+8b
0 = b^2+8b-180
0 = (b+18)(b-10)
b = {-18, 10}
throw out the negative solution leaving:
b = 10 feet (length of base)